From GlueXWiki
Jump to: navigation, search

March 1, 2012 FDC meeting


  1. Production Construction Tracking (Dave)
    • Production status
    • Testing at Blue Crab (Lubomir)
  2. Oxygen contamination in the first package: monitoring results, possible explanation and options to reduce it FDC ELOG (Beni, Bill, Lubomir)
  3. Engineering update (Bill)
  4. Electronics update (Chris)
  5. Other


Participants: Simon, Nick, Eugene, Beni, Chris, Bill, Dave, and Lubomir.


- Dave: Chris will come tomorrow to install the elements on the next wire frame; we decided it will be #23 and will have #13 (with the glued HV caps!) at the top of the package in case we have to fix it. Dave will do the electroplating tomorrow, as well. So, the wire frame and the cathodes for the next cell will be ready by Monday. Working on the rest of the cathodes for package #3 and started working also on the first cathode for package #4.

- Lubomir: first two cells of package #3 were installed, flushed with gas, conditioned and tested without problems within one week (~3 days for each). Yesterday we installed the third cell and started flushing with gas. Today after applying HV we found high dark current between the field and sense wires. We had similar problems before and we knew it is related to the epoxy used to glue the wires. It turned out that this is the first frame on which Casey used a new epoxy batch, both resin and hardener. Same epoxy was used on another three wire frames that will go next in the package. We will stop using this batch till we figure out if this is a problem.

Oxygen contamination in the first package

- Beni explained what have been tried so far. The assumption is there is trapped air in the volumes between the cathodes and ground plane connected to the main volumes through small openings. The package was turned vertical with the gas supply on the bottom and the exhaust on the top. A pump was installed at the end of the exhaust line through a rotarmeter (to regulate the flow) and oxygen sensor. The Gas flow was increased up to ~450 ccpm. There was some improvement in the oxygen percentage: from about ~12000 volumes per minute (vpm) to ~8000 vpm within three days of keeping the chamber vertical. Still not sure exactly how to convert vpm to percentage but at the gas supply we have ~200vpm and with the prototype we had ~800vpm.

- Bill has another explanation: syphoning air through the manifold fittings or the gas inlet at the gas spacers. Bill also started using ANSYS to model these effects. He suggested to measure the oxygen after the supply manifold and in the gauge tube. After the meeting Beni measured both: ~200vpm at one of the output of the supply manifold, and about ~8000vpm at the gauge tube. Therefore if there's syphoning effect it can be at the supply gas inlet in the spacer ring.

- Eugene estimated: ~10^5 cm^3 chamber volume, if half is trapped air the oxygen in the chamber will be 10^4 cm^3. At 200 cm^3/min flow and 1% oxygen in the exhaust will have 2 cm^3/min oxygen. so it will take ~100 hours to get the oxygen out. Since the oxygen percentage will drop it will take more.

- Bill suggested using fan pump instead of vacuum one to increase the flow through the chamber. With proper control one can keep the pressure in the chamber independent from the gas flow. Bill and Dave will investigate this option.


- Bill got permission to do pressure tests of the cooling system in the Hall. Will use first water then maybe fluorinert.

- Mark borrowed a moisture sensor for a couple of weeks that we will install at the exhaust.


- Grounding: Bill came with an idea how to make the common ground connections at the 6 points around the package with low material: using plastic clamps with Cu tape in between.